OpenStudy (anonymous):

Write the trigonometric expression in terms of sine and cosine, and then simplify. Tan(degree)/ Sec(degree)- Cos(degree)

OpenStudy (anonymous):

1: cos x 2. sin x 3. cot x 4. 1

OpenStudy (anonymous):

I don't understand what you wrote.

OpenStudy (anonymous):

Ok, I simplify it! Cos is the degree sec is degree tan degree! Each are = to x

OpenStudy (anonymous):

What would the answer be? I'm so confused, I'm sorry.

OpenStudy (anonymous):

Oh no your fine I understand what you mean... The answer is. 1

OpenStudy (anonymous):

Tan(x) = Sin(x) / Cos(x) Sec(x) = 1/Cos(x) => Tan(x)/Sec(x) - Cos(x) = (Sin(x)/Cos(x))/(1/Cos(x) - Cos(x) => multiplying numerator & denominator by Cos(x) we get Tan(x)/Sec(x)-Cos(x) = Sin(x)/1 - cos(x) = sin(x) - cos(x) Assumption: Cos(x) != 0

OpenStudy (anonymous):

Wait I looked at this again, its eather 0 or 1.